# Discrete-Time Open Quantum Walks for Vertex Ranking in Graphs

### v2

@@ Abstract
This article presents a new quantum PageRank algorithm on graphs using discrete-time open quantum walks. Google’s PageRank is an essential algorithm for arranging the web pages on the World Wide Web in classical computation. From a broader perspective, it is also a fundamental measure for quantifying the importance of vertices in a network. Similarly, the new quantum PageRank also serves to quantify the significance of a network’s vertices. In this work, we extend the concept of discrete-time open quantum walk on arbitrary directed and undirected graphs by utilizing the Weyl operators as Kraus operators. This new model of quantum walk is useful for building up the quantum PageRank algorithm, discussed in this article. We compare the classical PageRank and the newly defined quantum PageRank for different types of complex networks, such as the scale-free network, Erdős-Rényi random network, Watts-Strogatz network, spatial network, Zachary Karate club network, GNC, GN, GNR networks, Barabási and Albert network, etc. In addition, we study the convergence of the quantum PageRank process and its dependency on the damping factor $\alpha$. We observe that this quantum pagerank procedure is faster than many other proposals reported in the literature.
@@ 1 Introduction
There are multiple attempts to develop the PageRank algorithm in the literature of quantum computing which we can precisely classify into two different approaches. In one of the initial works on quantum PageRank (qPageRank), an adiabatic quantum computing algorithm was demonstrated [9]. The other proposals replace the random walk of classical PageRank with a quantum walk. In quantum computing we quantize the idea of random walk on the graphs with quantum walk [10]. There are different proposals of quantum walk, such as the discrete-time quantum walk [11], continuous-time quantum walk [12], Szegedy’s quantum walk [13, 14], open quantum walk [15], and many others. The qPageRank based on the discrete-time quantum walk is proposed in [16]. In [17] and [18] qPageRanks based on continuous-time open quantum walk are discussed. Articles [19, 20, 21], and [22] study different aspects of qPageRank based on Szegedy’s quantum walk. Article [23] compares the classical and quantum PageRanks on different types of networks. A number of classical computing aspects of qPageRank are discussed in [24]. Article [25] utilizes graph Fourier transformation on directed graphs and the quantum HHL algorithm [26] to calculate the classical page rank. A qPageRank based on the stochastic quantum walk is proposed in [27]. All these articles include numerical experiments on various models of complex networks to establish the justification for the efficiency of the newly developed algorithms, such as the outerplanar hierarchical network [19, 23, 25], the scale-free network [19, 23, 21, 22, 25, 27], Erdős–Rényi random network [23, 25, 27], Watts-Strogatz network [27], spatial network [27], Zachary Karate club network [27], random_k_out_graph [18], binary tree graph [21], GNC network [21], Barabási and Albert network [25], etc. The PageRank is an important measure of centrality on graphs. There are articles which quantize the notion of centrality using quantum walk for instance [28, 29]. The experimental realization of qPageRank and centrality measures are discussed in [30, 31, 32]. Although there are many proposals of quantum PageRank in literature, none of them exactly match with the classical PageRank for all graphs due to the difference between classical and quantum random walk models. As qPageRank is not unique, we assume that they should satisfy the following assumptions which were proposed in [16]:
@@
    The classical PageRank is defined on the directed graphs. Hence, the qPageRank should also be defined on the directed graphs.
@@
The classical algorithm for PageRank is simply an eigenvalue finding problem and does not consider quantum interference at all. Therefore, applying it to quantum networks is likely to yield an incorrect result. The literature crystallizes the crucial role of various quantum walks in developing the concept of qPageRank. Most of the quantum walks are originally defined on undirected graphs. To break this limitation, we introduce the open quantum walk, which is defined on directed graphs. Recall that the WWW is itself a directed graph, and the original idea of PageRank is defined on the directed graphs. Since we aim to define qPageRank for all graphs, it is crucial to properly define the quantum walk for all graphs, particularly for the directed graphs. The idea of an open quantum walk adds another dimension. Unlike the other quantum walks, the dynamics of the open quantum walk is non-unitary. It allows us to incorporate Markovian dynamics. It assures convergence of the instantaneous qPagerank sequence. For the other qPageRank algorithms, we require a sequence of time averages for convergence, but our particular choice of quantum walk eliminates this requirement. Our article uses the discrete-time open quantum walk, which was not employed before to quantify PageRank because it was mostly talked about on a few undirected regular graphs, like the infinite path graph or $\mathbb{Z}$, cycle [15, 34, 35]. To construct qPageRank with this quantum walk, we extend its definition to all graphs. This quantum walk model primarily employs the Kraus operators . We observe that the Weyl operators [36, 37] can also be utilized to construct the Kraus operators [38]. We extend the idea of an open quantum walk for arbitrary directed and undirected graphs. To the best of our knowledge, we employ the Weyl operators to define the open quantum walk for the first time in literature.
@@
This article is distributed as follows: Section 2 compiles all the preliminary ideas for this article. It has three subsections. We begin with an introduction to graph theory, which is the mathematical foundation of the network analysis. In section 2.2, we discuss the original idea of classical PageRank. We end this section with a discussion on quantum states and channels in higher dimensions. In Section 3, we introduce a new concept by demonstrating a generalization of the discrete-times open quantum walk for any directed graphs. Section 4 models how an internet surfer can restart surfing from another randomly chosen node in a quantum mechanical way, which is a crucial assumption in the classical PageRank algorithm. The qPageRank of a graph’s vertices is the focus of Section 5. We discuss our qPageRank algorithm in Procedure 1. In this section, we justify that a classical computer can calculate the new PageRank algorithm in polynomial time on an $n$-dimensional Hilbert space. We divide our numerical experiments into four subsections in Section 6. In the first two subsections, we compare the classical and quantum PageRanks for a number of undirected and directed graphs, respectively. We dedicate the third subsection to the convergence of the qPageRank procedure. The last subsection presents the dependency of qPageRank on the primitivity parameter. Then, we conclude this article.
@@ 2.1 An introduction to graph theory
A graph $G=(V(G),E(G))$ is a combinatorial object consisting of a set of vertices $V(G)$ and a set of edges $E(G)\subset V(G)\times V(G)$ [39]. Throughout this article, $n$ denotes the number of vertices in a graph $G$. An edge $(u,v)$ between the vertices $u$ and $v$ is said to be incident to the vertices $u$ and $v$. Also, $u$ and $v$ are said to be the neighbors of each others. We employ different types of edges in this article. We denote a directed edge or link from the vertex $u$ to a vertex $v$ by $\overrightarrow{(u,v)}$. An undirected edge between vertices $u$ and $v$ is denoted by $(u,v)$, which can be considered as a combination of two directed edges, one from $u$ to $v$, that is $\overrightarrow{(u,v)}$ and another from $v$ to $u$, that is $\overrightarrow{(v,u)}$. In this article, the graphs have no loop on the vertices, which is an edge to join a vertex with itself. Also, there are no weights on the vertices and on the edges. All the edges of a directed graph are directed. Also, all the edges of an undirected graph are undirected, throughout this article.
@@
For our numerical experiments we use a number of graphs and networks, which may be directed or undirected. For instance, we use the path graph, cycle graph, complete graph, star graph, balanced tree graph, wheel graph, etc. Also we employ the scale-free network, Erdős–Rényi random network, Watts-Strogatz network, Zachary Karate club network, random_k_out_graph, tree graph, GNC, GN, GNR network, Barabási and Albert network, etc. We avoid individual descriptions of them to make the article shorter. The reader may go through the documentation of NetworkX for the further details [40].
@@ 2.2 An introduction to Google’s PageRank
To build up the mathematical model of information retrieval from WWW, we utilize the idea of an internet graph, which is a directed graph. Recall that the vertices of the internet graph represent webpages in WWW. The directed edges from one vertex to another represent the hyperlinks from one page to another, correspondingly. Although, Google’s PageRank was originally developed and utilized to quantify the importance of webpages, it is a crucial measure to quantify the importance of vertices in any network.
@@
Initially, $r_{0}(v)=\frac{1}{n}$ for all vertices $v$, where $n$ is the total number of vertices in $\overrightarrow{G}$. A PageRank vector $\Pi^{(k)}$ is a vector of order $1\times n$ which contains the PageRanks of the vertices at $k$-th iteration for $k=0,1,2,\dots$. Initially, we consider that the walker starts walking from any vertex of the graph, which is chosen uniformly and randomly. Therefore, the initial page rank vector is $\Pi^{(0)}=\frac{1}{n}(1,1,\dots 1(n$-times)). Here, the reader should notice that $\Pi^{(k)}$ is not a quantum state vector, as it may not be a normalized vector. In the next subsection, we describe quantum states in detail. The hyperlink matrix of the graph is defined by $H=(h_{u,v})_{n\times n}$, where
@@
$\Pi^{(k+1)}=\Pi^{(k)}H.$ (4)
for $k=0,1,2,\dots$. We want to convert this iterative process to a Markov process [41] on graphs. The matrix $H$ depends on $\overrightarrow{G}$. It is not always a stochastic matrix. Therefore, the convergence of this iterative process is not guaranteed. To ensure convergence we make the following adjustments to the matrix $H$:
@@
    Stochasticity adjustment: Let $v$ be a dangling node. Recall that there are no outgoing edges from $v$ to any other vertex. After entering a dangling node, we assume that the walker can move to any page randomly. Therefore, we replace the matrix $H$ by $S=H+\frac{1}{n}\ket{a}\bra{e}$ |, where |$\ket{a}$ and |$\ket{e}$ are the column vectors of order $n$. The vector |$\ket{e}$ is the all $1$ vector. The vector |$\ket{a}$ consists of the elements $a_{v}=1$ if $v$ is a dangling node and 00 otherwise. We can prove that the matrix $S$ is a stochastic matrix.
@@
    $\mathcal{G}=\alpha S+\frac{(1-\alpha)}{n}ee^{T}.$ (5)
@@
It can be proved that, the matrix $\mathcal{G}$ is a primitive matrix, which is both irreducible and aperiodic [42]. This crucial characteristic of $\mathcal{G}$ ensures that $\lim_{k\rightarrow\infty}\Pi^{(k)}=\Pi$ exists. We call $\Pi$ as a PageRank vector which is unique for a graph.
@@ 2.3 Quantum states and channels in higher dimensional system
In quantum mechanics and information theory, a pure state is equivalent to a normalized column vector in a Hilbert space $\mathcal{H}^{n}$ of dimension $n$. In this article, we utilize Dirac’s bra-ket notation, only to denote the state vectors. Therefore, |$\ket{\psi}$ denotes a pure state and $\bra{\psi}$ | is its conjugate transpose. There are a number of vectors in this article, such as the PageRank vector, which may not satisfy the condition of normalization. We do not use the bra-ket notation for them. We consider the standard basic of $\mathcal{H}^{n}$ which is {|$\{\ket{i}:\ket{i}=(0,0,\dots 1(i\text{-th position}),\dots 0)^{\dagger}\}$. A mixed quantum state in $\mathcal{H}^{n}$ is represented by a density matrix $\rho$, which is a positive semidefinite Hermitian matrix of order $n$ and $\operatorname{trace}(\rho)=1$.
@@ 3 Discrete-time open quantum walk on arbitrary directed graph
$\begin{split}\rho^{(t+1)}_{K}&=\Psi_{K}(\rho^{(t)})=\sum_{v\in V(G)}\sum_{u\in\operatorname{Nbd}(v)\cup\{v\}}K_{v,u}\rho_{v}^{(t)}\otimes\ket{v}\bra{v}K_{v,u}^{\dagger}\\ &=\sum_{v\in V(G)}\sum_{u\in\operatorname{Nbd}(v)\cup\{v\}}C_{v,u}\rho_{v}^{(t)}C_{v,u}^{\dagger}\otimes S_{v,u}\ket{v}\bra{v}S_{v,u}^{\dagger}\\ &=\sum_{v\in V(G)}\sum_{u\in\operatorname{Nbd}(v)\cup\{v\}}C_{v,u}\rho_{v}^{(t)}C_{v,u}^{\dagger}\otimes\ket{u}\braket{v}{v}\braket{v}{v}\bra{u}\\ &=\sum_{v\in V(G)}\sum_{u\in\operatorname{Nbd}(v)\cup\{v\}}C_{v,u}\rho_{v}^{(t)}C_{v,u}^{\dagger}\otimes\ket{u}\bra{u}.\end{split}$ (19)
@@
$\begin{split}\rho^{(t+1)}_{K}&=\sum_{v\in V(G)}\sum_{u\in\operatorname{Nbd}(v)\cup\{v\}}C_{v,u}\rho_{v}^{(t)}C_{v,u}^{\dagger}\otimes\ket{u}\bra{u}\\ &=\sum_{u\in V(G)}\sum_{v\in B(u)\cup\{u\}}C_{v,u}\rho_{v}^{(t)}C_{v,u}^{\dagger}\otimes\ket{u}\bra{u}.\end{split}$ (20)
@@
The equation (21) reduces the size of Hilbert space in our consideration. The state |$\ket{u}\bra{u}$ | in equation (17) remains unchanged in all the steps. We update the coin space corresponding to a vertex from time to time using equation (21). Therefore, the Hilbert space for our calculations is $\mathcal{H}^{n}$, which is a space of dimension $n$.

### v3

@@ Abstract
This article presents a new quantum PageRank algorithm on graphs using discrete-time open quantum walks. Google’s PageRank is a widely used algorithm for ranking the web pages on the World Wide Web in classical computation. From a broader perspective, it is also a fundamental measure for quantifying the importance of vertices in a network. Similarly, the new quantum PageRank also serves to quantify the significance of a network’s vertices. In this work, we extend the concept of discrete-time open quantum walk on arbitrary directed and undirected graphs by utilizing the Weyl operators as Kraus operators. This new model of quantum walk is useful for building up the quantum PageRank algorithm, discussed in this article. We compare the classical PageRank and the newly defined quantum PageRank for different types of complex networks, such as the scale-free network, Erdős-Rényi random network, Watts-Strogatz network, spatial network, Zachary Karate club network, GNC, GN, GNR networks, Barabási and Albert network, etc. In addition, we study the convergence of the quantum PageRank process and its dependency on the damping factor $\alpha$. We observe that this quantum PageRank procedure is faster than many other proposals reported in the literature.
@@ 1 Introduction
There are multiple attempts to develop the PageRank algorithm in the literature of quantum computing which we can precisely classify into two different approaches. In one of the initial works on quantum PageRank (qPageRank), an adiabatic quantum computing algorithm was demonstrated [9]. The other proposals replace the random walk of classical PageRank with a quantum walk. In quantum computing we quantize the idea of random walk on the graphs with quantum walk [10]. There are different proposals of quantum walk, such as the discrete-time quantum walk [11, 12, 13], continuous-time quantum walk [14, 15], open quantum walk [16], and many others. A number of articles that use continuous-time quantum walk for developing qPageRank are [17, 18, 19, 20]. The qPageRank based on the discrete-time quantum walk is proposed in [21]. There are articles which quantize the notion of centrality using quantum walk for instance [18, 19]. In [22] qPageRanks based on continuous-time open quantum walk are discussed. Articles [23, 24, 25], and [26] study different aspects of qPageRank based on Szegedy’s quantum walk. Article [27] compares the classical and quantum PageRanks on different types of networks. A number of classical computing aspects of qPageRank are discussed in [28]. Article [29] utilizes graph Fourier transformation on directed graphs and the quantum HHL algorithm [30] to calculate the classical PageRank. A qPageRank based on the stochastic quantum walk is proposed in [31]. All these articles include numerical experiments on various models of complex networks to establish the justification for the efficiency of the newly developed algorithms, such as the outerplanar hierarchical network [23, 27, 29], the scale-free network [23, 27, 25, 26, 29, 31], Erdős–Rényi random network [27, 29, 31], Watts-Strogatz network [31], spatial network [31], Zachary Karate club network [31], random_k_out_graph [17], binary tree graph [25], GNC network [25], Barabási and Albert network [29], etc. The PageRank is an important measure of centrality on graphs. The experimental realization of qPageRank and centrality measures are discussed in [32, 20, 33]. Although there are many proposals of quantum PageRank in literature, none of them exactly match with the classical PageRank for all graphs due to the difference between classical and quantum random walk models. As qPageRank is not unique, we assume that they should satisfy the following assumptions which were proposed in [21]:
@@
    The classical PageRank is defined on directed graphs. Hence, the qPageRank should also be defined on directed graphs.
@@
The classical algorithm for PageRank is an eigenvalue finding problem and does not consider quantum interference at all. Therefore, it is essential to develop a quantum mechanical algorithm for ranking the vertices in a network. The literature crystallizes the crucial role of various quantum walks in developing the concept of qPageRank. Most of the quantum walks are originally defined on undirected graphs. To break this limitation, we introduce the open quantum walk, which is defined on directed graphs. Recall that the WWW is itself a directed graph, and the original idea of PageRank is defined on directed graphs. Since we aim to define qPageRank for all graphs, it is crucial to properly define the quantum walk for all graphs, particularly for directed graphs. The idea of an open quantum walk adds another dimension. Unlike the other quantum walks, the dynamics of the open quantum walk is non-unitary. It allows us to incorporate Markovian dynamics. It assures convergence of the instantaneous qPagerank sequence. For the other qPageRank algorithms, we require a sequence of time averages for convergence, but our particular choice of quantum walk eliminates this requirement. Our article uses the discrete-time open quantum walk, which was not employed before to quantify PageRank because it was mostly studied on a few undirected regular graphs, like the infinite path graph. or $\mathbb{Z}$, cycle [16, 35, 36]. To construct qPageRank with this quantum walk, we extend its definition to all graphs. This quantum walk model primarily employs the Kraus operators . We observe that the Weyl operators [37, 38] can also be utilized to construct the Kraus operators [39]. We extend the idea of an open quantum walk for arbitrary directed and undirected graphs. To the best of our knowledge, we employ the Weyl operators to define the open quantum walk for the first time in literature.
@@
This article is organized as follows: Section 2 compiles all the preliminary ideas for this article. It has three subsections. We begin with an introduction to graph theory, which is the mathematical foundation of the network analysis. In section 2.2, we discuss the original idea of classical PageRank. We end this section with a discussion on quantum states and channels in higher dimensions. In Section 3, we introduce a new concept by demonstrating a generalization of the discrete-times open quantum walk for any directed graphs. Section 4 models how an internet surfer can restart surfing from another randomly chosen node in a quantum mechanical way, which is a crucial assumption in the classical PageRank algorithm. The qPageRank of a graph’s vertices is the focus of Section 5. We discuss our qPageRank algorithm in Procedure 1. In this section, we justify that a classical computer can calculate the new PageRank algorithm in polynomial time on an $n$-dimensional Hilbert space. We divide our numerical experiments into four subsections in Section 6. In the first two subsections, we compare the classical and quantum PageRanks for a number of undirected and directed graphs, respectively. We dedicate the third subsection to the convergence of the qPageRank procedure. The last subsection presents the dependency of qPageRank on the primitivity parameter. Then, we conclude this article.
@@ 2.1 An introduction to graph theory
A graph $G=(V(G),E(G))$ is a combinatorial object consisting of a set of vertices $V(G)$ and a set of edges $E(G)\subset V(G)\times V(G)$ [40]. Throughout this article, $n$ denotes the number of vertices in a graph $G$. An edge $(u,v)$ between the vertices $u$ and $v$ is said to be incident to the vertices $u$ and $v$. Also, $u$ and $v$ are said to be the neighbors of each other. We employ different types of edges in this article. We denote a directed edge or link from the vertex $u$ to a vertex $v$ by $\overrightarrow{(u,v)}$. An undirected edge between vertices $u$ and $v$ is denoted by $(u,v)$, which can be considered as a combination of two directed edges, one from $u$ to $v$, that is $\overrightarrow{(u,v)}$ and another from $v$ to $u$, that is $\overrightarrow{(v,u)}$. In this article, the graphs have no loop on the vertices, which is an edge to join a vertex with itself. Also, there are no weights on the vertices and on the edges. All the edges of a directed graph are directed. Also, all the edges of an undirected graph are undirected, throughout this article.
@@
For our numerical experiments we use a number of graphs and networks, which may be directed or undirected. For instance, we use the path graph, cycle graph, complete graph, star graph, balanced tree graph, wheel graph, etc. Also we employ the scale-free network, Erdős–Rényi random network, Watts-Strogatz network, Zachary Karate club network, random_k_out_graph, tree graph, GNC, GN, GNR network, Barabási and Albert network, etc. We avoid individual descriptions of them to make the article shorter. The reader may go through the documentation of NetworkX for further details [41].
@@ 2.2 An introduction to Google’s PageRank
To build up the mathematical model of information retrieval from WWW, we utilize the idea of an internet graph, which is a directed graph. Recall that the vertices of the internet graph represent webpages in WWW. The directed edges from one vertex to another represent the hyperlinks from one page to another, correspondingly.
@@
Initially, $r_{0}(v)=\frac{1}{n}$ for all vertices $v$, where $n$ is the total number of vertices in $\overrightarrow{G}$. Let $\Pi^{(k)}$ be a vector of order $1\times n$ which contains the PageRanks of the vertices at $k$-th iteration for $k=0,1,2,\dots$. Initially, we consider that the walker starts walking from any vertex of the graph, which is chosen uniformly and randomly. Therefore, the initial PageRank vector is $\Pi^{(0)}=\frac{1}{n}(1,1,\dots 1(n$-times)). Here, the reader may notice that $\Pi^{(k)}$ represents a probability distribution for all $k$. The hyperlink matrix of the graph is defined by $H=(h_{u,v})_{n\times n}$, where
@@
$\Pi^{(k+1)}=\Pi^{(k)}H,$ (4)
for $k=0,1,2,\dots$. To ensure convergence of this iterative process we make the following adjustments to the matrix $H$:
@@
    Stochasticity adjustment: Let $v$ be a dangling node. Recall that there are no outgoing edges from $v$ to any other vertex. After entering a dangling node, we assume that the walker can move to any page randomly. Therefore, we replace the matrix $H$ by $S=H+\frac{1}{n}ae^{\dagger}$, where $a$ and $e$ are the column vectors of order $n$. The vector $e$ is the all $1$ vector. The vector $a$ consists of the elements $a_{v}=1$ if $v$ is a dangling node and 00 otherwise.
@@
    $\mathcal{G}=\alpha S+\frac{(1-\alpha)}{n}ee^{\dagger}.$ (5)
@@
It can be proved that $\mathcal{G}$ ensures the convergence $\{\Pi^{(k)}\}$. Therefore, $\lim_{k\rightarrow\infty}\Pi^{(k)}=\Pi$ exists. We call $\Pi$ as a PageRank vector which is unique for a graph.
@@ 2.3 Quantum states and channels in higher dimensional system
In quantum mechanics and information theory, a pure state is equivalent to a normalized column vector in a Hilbert space $\mathcal{H}^{n}$ of dimension $n$. In this article, we utilize Dirac’s bra-ket notation, only to denote the state vectors. Therefore, |$\ket{\psi}$ denotes a pure state and $\bra{\psi}$ | is its conjugate transpose. There are a number of vectors in this article, such as the PageRank vector, which may not satisfy the condition of normalization. We do not use the bra-ket notation for them. We consider the standard basis of $\mathcal{H}^{n}$ which is {|$\{\ket{i}:\ket{i}=(0,0,\dots 1(i\text{-th position}),\dots 0)^{\dagger}\}$. A mixed quantum state in $\mathcal{H}^{n}$ is represented by a density matrix $\rho$, which is a positive semidefinite Hermitian matrix of order $n$ and $\operatorname{trace}(\rho)=1$.
@@ 3 Discrete-time open quantum walk on arbitrary directed graph
$\begin{split}\rho^{(t+1)}&=\Psi_{K}(\rho^{(t)})=\sum_{v\in V(G)}\sum_{u\in\operatorname{Nbd}(v)\cup\{v\}}K_{v,u}\rho_{v}^{(t)}\otimes\ket{v}\bra{v}K_{v,u}^{\dagger}\\ &=\sum_{v\in V(G)}\sum_{u\in\operatorname{Nbd}(v)\cup\{v\}}C_{v,u}\rho_{v}^{(t)}C_{v,u}^{\dagger}\otimes S_{v,u}\ket{v}\bra{v}S_{v,u}^{\dagger}\\ &=\sum_{v\in V(G)}\sum_{u\in\operatorname{Nbd}(v)\cup\{v\}}C_{v,u}\rho_{v}^{(t)}C_{v,u}^{\dagger}\otimes\ket{u}\braket{v}{v}\braket{v}{v}\bra{u}\\ &=\sum_{v\in V(G)}\sum_{u\in\operatorname{Nbd}(v)\cup\{v\}}C_{v,u}\rho_{v}^{(t)}C_{v,u}^{\dagger}\otimes\ket{u}\bra{u}.\end{split}$ (19)
@@
$\begin{split}\rho^{(t+1)}&=\sum_{v\in V(G)}\sum_{u\in\operatorname{Nbd}(v)\cup\{v\}}C_{v,u}\rho_{v}^{(t)}C_{v,u}^{\dagger}\otimes\ket{u}\bra{u}\\ &=\sum_{u\in V(G)}\sum_{v\in B(u)\cup\{u\}}C_{v,u}\rho_{v}^{(t)}C_{v,u}^{\dagger}\otimes\ket{u}\bra{u}\\ &=\sum_{v\in V(G)}\sum_{u\in B(v)\cup\{v\}}C_{u,v}\rho_{u}^{(t)}C_{u,v}^{\dagger}\otimes\ket{v}\bra{v},~{}\text{by interchanging the indices $u$ and $v$.}\end{split}$ (20)
@@
The equation (21) reduces the size of Hilbert space in our consideration. In general, the dynamics of our quantum walk formally takes place in a Hilbert space of dimension $n^{2}$. But, because of the separable Kronecker product structure of the states and operators involved, one can restrict the actual computations to a Hilbert subspace of dimension $n$. This is a known behavior of discrete-time quantum walks on graphs.[21, 23, 24]. The state |$\ket{v}\bra{v}$ | in equation (17) remains unchanged in all the steps. We update the coin space corresponding to a vertex from time to time using equation (21). Therefore, the Hilbert space for our calculations is $\mathcal{H}^{n}$, which is a space of dimension $n$.